# EmbodiedAgent: A Scalable Hierarchical Approach to Overcome Practical Challenge in Multi-Robot Control

### v1

@@ I INTRODUCTION
In response to these limitations, this work introduces a hierarchical Embodied system with an Agent-based planner, named EmbodiedAgent. EmbodiedAgent leverages a next-action prediction paradigm to establish a heterogeneous multi-robot control system. The core agent generates a single action and its corresponding arguments per inference, terminating upon receiving an end-of-planning signal, thus ensuring a controlled and concise execution process. To address the aforementioned challenges, we enhance the planner’s robustness and generalizability through supervised fine-tuning. We present MultiPlan+, a large-scale dataset comprising 100 scenarios with over 18,000 tasks, enriched with a subset of impractical cases to mitigate hallucinations. Additionally, we develop an agent based on a fine-tuned language model equipped with function calling capabilities and structured memory. Specifically, robot skills, termination signals, and error signals related to impractical cases are encapsulated as tools, while planning history is organized within the structured memory. For low-level execution, we employ specialized policies trained on individual basic tasks to ensure reliable and robust performance. Furthermore, we propose a comprehensive Robot Planning Assessment Schema (RPAS), grounded in MRED [4], which moves beyond error-type diagnostics to emphasize stratified success rates assessed through both human evaluation and automated grading. Code and dataset are open-sourced  [^1] https://github.com/HaronW/EmbodiedAgent. In summary, our main contributions are as follows:
@@ II PROBLEM FORMULATION
Given the planning state $s_{t}=\{\mathbf{M},\mathbf{E},\mathbf{P}_{t}\}$ at planning step $t$, the objective of the agent-based planner $\pi$ is to generate the next planning $p_{t+1}$. The next action prediction planning loop of EmbodiedAgent ends with the termination signal $\epsilon_{end}$ or throwing impractical errors, formulated as: =$\mathcal{E}=\{\epsilon_{LoA},\epsilon_{LoO},\epsilon_{LoS},\epsilon_{LoL}\}$. Here, $\epsilon_{end}$ signifies successful task completion, while the other signals denote various execution failures. The complete planning objective can be formalized as finding an optimal action sequence that successfully accomplishes the mission while satisfying all operational constraints:
[…]
@@ III-A EmbodiedAgent Architecture
4:while step $<$ threshold and not termination do
@@
6:     $res\leftarrow\pi(Q)$triangleright get LLM response
@@ III-B MultiPlan+ Dataset
MultiPlan+ is a large-scale dataset for heterogeneous multi-robot planning, extended from MultiPlan. The dataset consists of over 18,000 data entries derived from more than 3,400 unique tasks within 100 diverse scenarios, including office, domestic, urban street area, exploratory environments, and etc. The distribution of the data samples are shown in fig. 3. Each data entry is annotated with task specification, environmental states, planning memory, and ground-truth action responses. The environment description details the workspace with position points, robot team configurations, objects, and involved users.
@@ III-C3 Multi-Robot Planning Error Diagnosis (MRED)
Model Infer type $ASR_{top-k}$ Expert grading UE PoE PlE SE EE RPAS
@@ IV-B Multi-Robot System for Office Service
To validate the robustness of EmbodiedAgent in real-world settings, we designed an office service scenario for testing. The mission description, environment setups, planning sequence, and a conceptual illustration are shown in fig. 4. For low-level execution, we employ the Action Chunking Transformer (ACT) to train the robot arm on individual tasks and utilize the SDK for the robot dog. To enable autonomous task execution by the robot arm, we first collected demonstration data through teleoperation, where human operators remotely controlled the arm. A total of 50 demonstration episodes were collected for each task, with each episode lasting approximately 15-20 seconds. Video data was captured from three camera views: a wrist camera view, a top-down view, and a view of the basket mounted on the back of the robot dog. These diverse viewpoints provided comprehensive observational data, which were then used to train a policy using the ACT algorithm. On the other hand, the robot utilize depth camera for navigation and collision avoidance. This experiment assessed the effectiveness of the robot arm and robot dog system in performing complex, coordinated tasks in dynamic and realistic office settings. It evaluated the system’s ability to transfer learned behaviors to new, unseen office scenarios, focusing on adaptability, efficiency, and overall performance under real-world conditions.

### v2

@@ I INTRODUCTION
In response to these limitations, this work introduces a hierarchical Embodied system with an Agent-based planner, named EmbodiedAgent. EmbodiedAgent leverages a next-action prediction paradigm to establish a heterogeneous multi-robot control system. The core agent generates a single action and its corresponding arguments per inference, terminating upon receiving an end-of-planning signal, thus ensuring a controlled and concise execution process. To address the aforementioned challenges, we enhance the planner’s robustness and generalizability through supervised fine-tuning. Extended from previous work MultiPlan [4], we present MultiPlan+, a large-scale dataset comprising 100 scenarios with over 18,000 tasks, enriched with a subset of impractical cases to mitigate hallucinations. Additionally, we develop an agent based on a fine-tuned language model equipped with function calling capabilities and structured memory. Specifically, robot skills, termination signals, and error signals related to impractical cases are encapsulated as tools, while planning history is organized within the structured memory. For low-level execution, we employ specialized policies trained on individual basic tasks to ensure reliable and robust performance. Furthermore, we propose a comprehensive Robot Planning Assessment Schema (RPAS), which moves beyond error-type diagnostics to emphasize stratified success rates assessed through both human evaluation and automated grading. Code and dataset are open-sourced  [^1] https://github.com/HaronW/EmbodiedAgent. In summary, our main contributions are as follows:
@@ II PROBLEM FORMULATION
Given the planning state $s_{t}=\{\mathbf{M},\mathbf{E},\mathbf{P}_{t}\}$ at planning step $t$, the objective of the agent-based planner $\pi$ is to generate the next planning $p_{t+1}$. The next action prediction planning loop of EmbodiedAgent ends with the termination signal $\epsilon_{end}$ or throwing impractical errors, formulated as: $\mathcal{E}=\{\epsilon_{LoA},\epsilon_{LoO},\epsilon_{LoS},\epsilon_{LoL}\}$. Here, $\epsilon_{end}$ signifies successful task completion, while the other signals denote various execution failures. The complete planning objective can be formalized as finding an optimal action sequence that successfully accomplishes the mission while satisfying all operational constraints:
$\begin{gathered}p_{t+1}=\mathop{\pi}(\mathbf{P}_{t}\mid\mathbf{M},\mathbf{E}),\\ p_{t+1}\in\{a,\epsilon_{end},\mathcal{E}\},\end{gathered}$ (1)
@@ III-A EmbodiedAgent Architecture
4:while step << threshold and not termination do
@@
6:  $res\leftarrow\pi(Q)$ $\triangleright$ get LLM response
@@ III-B MultiPlan+ Dataset
MultiPlan+ is a large-scale dataset for heterogeneous multi-robot planning, extended from MultiPlan. The dataset consists of over 18,000 data entries derived from more than 3,400 unique tasks within 100 diverse scenarios, including office, domestic, urban street area, exploratory environments, and etc. The distribution of the data samples are shown in fig. 3. Each data entry is annotated with task specification, environmental states, planning memory, and ground-truth action responses. The environment description details the workspace with position points, robot team configurations, objects, and involved users. Different from the MultiPlan dataset, MultiPlan+ adopts the next action prediction paradigm and has been reviewed for diversity.
@@ III-C3 Multi-Robot Planning Error Diagnosis (MRED)
Model Infer type $ASR_{top-k}$ Expert grading UE PoE PlE SE $E$ RPAS
@@ IV-B Multi-Robot System for Office Service
To validate the robustness of EmbodiedAgent in real-world settings, we designed an office service scenario for testing. The mission description, environment setups, planning sequence, and a conceptual illustration are shown in fig. 4. For low-level execution, we employ the Action Chunking Transformer (ACT) to train the robot arm on individual tasks and utilize the SDK for the robot dog. To enable autonomous task execution by the robot arm, we first collected demonstration data through teleoperation, where human operators remotely controlled the arm. A total of 50 demonstration episodes were collected for each task, with each episode lasting approximately 15-20 seconds. Video data was captured from three camera views: a wrist camera view, a top-down view, and a view of the basket mounted on the back of the robot dog, as shown in fig. 5a. This experiment assessed the effectiveness of the robot arm and robot dog system in performing complex, captured frames are shown in fig. 5b. It evaluated the system’s ability to transfer learned behaviors to new, unseen office scenarios, focusing on adaptability, efficiency, and overall performance under real-world conditions.